OpenStudy (anonymous):

what is x= 11 (Mod 14)?

OpenStudy (anonymous):

x = 14k+11 where k = 0, 1, 2, ...

OpenStudy (anonymous):

but what would the reduced form be?

OpenStudy (anonymous):

reduced form? i think that's the simplest way to express x

OpenStudy (anonymous):

i mean the original equation is 11 mod 7 so could i not write that as 4 mod 7

OpenStudy (anonymous):

x = 11 mod 7 is equivalent to x = 4 mod 7 and x=7k+4 k = 0, 1, 2, ...

OpenStudy (anonymous):

where did the 14 come in?

OpenStudy (anonymous):

because the original problem was to find the smallest specific solution for x=1 (mod 2) and x=4 (mod 7)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

so what im gonna do is reduce x= 11 (mod 7) into x =4 (mod 7) and then just make it x=4 (mod 14)

OpenStudy (anonymous):

no... tha won't satisfy the original. think about it... whatever x is, it has to be odd. otherwise, x = 0 mod 2. therefore, it would be x = 7k + 4, k = 1, 3, 5, ... or the other way to write it would be x = 14n + 11, n = 0, 1, 2, ... you'll see you get the same numbers... k = 1 is the same as n = 0, k = 3 is the same as n = 1, and so on.

OpenStudy (anonymous):

18 = 4 mod 14 but 18 = 0 mod 2

OpenStudy (anonymous):

so i should just leave it as x=11 (mod 7) right

OpenStudy (anonymous):

i mean x= 11 mod 14

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wolfram alpha says 1(mod 2) =11 (mod 14) is false

OpenStudy (anonymous):

it should be a solution but its not working

OpenStudy (loser66):

@wednesday09876 you said that the original problem is x=1 (mod 2) x =4 (mod 7) Q: find the smallest number of x?

OpenStudy (anonymous):

Question: Give the smallest specific solution and the modulus for its congruence class.

OpenStudy (anonymous):

x = 11. that is the smallest number that satisfies x = 1 mod 2 and x = 4 mod 7

OpenStudy (anonymous):

okay. so my final answer is going to be x=11 (mod 14) right?

OpenStudy (anonymous):

1 mod 2 <> 11 mod 14... only = in certain cases.

OpenStudy (anonymous):

if they want the smallest number, it's 11. if they want all numbers then it's x = 11 mod 14,

OpenStudy (anonymous):

great thanks.

OpenStudy (anonymous):

you're welcome